OpenStudy (anonymous):

Determine the number and type of solutions for x2 + 4x + 3=0.

OpenStudy (anonymous):

2 Rational 1 Rational Double 2 Irrational 2 Complex

OpenStudy (anonymous):

Discriminat= 16-12 = 4

jimthompson5910 (jim_thompson5910):

D=b^2-4ac D=(4)^2-4(1)(3) D=16-4(1)(3) D=16-12 D=4 So the discriminant is D=4 Since the discriminant is greater than zero, this means that there are two real solutions. Since the discriminant is a perfect square, this means that these two real solutions are also rational. So the answer is choice A

OpenStudy (anonymous):

2 Rational.

jhonyy9 (jhonyy9):

(x+1)(x+3)

jhonyy9 (jhonyy9):

solutions are -1 and -3